# Convolution estimates for the Boltzmann gain operator with hard-spheres

### v3

@@ 4.2. The $L_{v}^{\infty}$– estimate for $Q_{1}$
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{p_{2}}|g_{k}(v_{1}^{*})|^{lp_{2}}||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lp_{2}}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Since $l>2$, (2.26) implies $I_{2}\lesssim{\langle}v{\rangle}^{p_{2}-2}\|g_{k}\|_{L_{v}^{p_{2}}}^{p_{2}}$.
In order to estimate $I_{1}$, we use the definition (4.15) of $\mathcal{B}_{1}(v)$, followed by (2.6), (2.12), (2.13) and the substitution $y:=u=v-v_{1}$ to write
$I_{1}(v)$ $\leq\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(u))|^{q_{2}}\mathds{1}_{|v|>10}\frac{\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{q_{2}/p_{2}}}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(y))|^{q_{2}}\mathds{1}_{|v|>10}\frac{\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{q_{2}/p_{2}}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ on the right-hand side of the above bound, to obtain
$I_{1}(v)$ $\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q_{2}/p_{2}}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q_{2}/p_{2}}}\,d\sigma\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{0}^{\pi/2}\frac{\mathds{1}_{\sin\theta\geq{\langle}v{\rangle}^{-1/2}}\chi(2\cos^{2}\theta-1)}{\cos^{2}\theta(\sin\theta)^{\frac{2q_{2}}{p_{2}}-1}}\,d\theta\,d\nu$
$=\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{0}^{\pi/2}\frac{\mathds{1}_{\sin\theta\geq{\langle}v{\rangle}^{-1/2}}\chi(2\cos^{2}\theta-1)\cos\theta}{\cos^{3}\theta(\sin\theta)^{\frac{2q_{2}}{p_{2}}-1}}\,d\theta\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{0}^{\pi/4}\frac{\mathds{1}_{\sin\theta\geq{\langle}v{\rangle}^{-1/2}}\cos\theta}{(\sin\theta)^{\frac{2q_{2}}{p_{2}}-1}}\,d\theta\,d\nu$
$=\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{{\langle}v{\rangle}^{-1/2}}^{1/\sqrt{2}}x^{1-\frac{2q_{2}}{p_{2}}}\,dx\,d\nu$
$\lesssim{\langle}v{\rangle}^{\frac{q_{2}}{p_{2}}-1}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\,d\nu={\langle}v{\rangle}^{\frac{q_{2}}{p_{2}}-1}\|f_{l}\|_{L_{v}^{q_{2}}}^{q_{2}},$
where we used the substitution $x:=\sin\theta$, and the fact that $p_{2}<2<q_{2}$ to estimate the integral in $x$. Note that the right-hand side of the above estimate is growing as $|v|\to\infty$, however the decay of $I_{2}$ is sufficient to counteract this growth.
Indeed, combining (4.20) with the estimates for $I_{1},I_{2}$ and using the exponents’ condition, we obtain
(4.21) ${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim|v|^{\frac{1}{p_{2}}-\frac{1}{q_{2}}+1-\frac{2}{p_{2}}}\|f_{l}\|_{L_{v}^{q_{2}}}\|g_{k}\|_{L_{v}^{p_{2}}}=\|f_{l}\|_{L_{v}^{q_{2}}}\|g_{k}\|_{L_{v}^{p_{2}}},\quad 1<p_{2}<2.$
$\bullet$ Case $p_{2}=1$ : We use again (4.19) and the definition of $\mathcal{B}_{1}(v)$ followed by (2.26) to bound
@@
${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim\|f_{l}\|_{L_{v}^{\infty}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|g_{k}(v_{1}^{*})|}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|\geq{\langle}v{\rangle}^{-1/2}}\,d\sigma\,dv_{1}$
$\leq\|f_{l}\|_{L_{v}^{\infty}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{2}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Since $l>2$, (2.26) implies
(4.22) ${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|\lesssim\|f_{l}\|_{L_{v}^{\infty}}\|g_{k}\|_{L_{v}^{1}}.$
Combining (4.21), (4.22), we conclude
(4.23) $\|{\langle}v{\rangle}^{k}Q_{1}^{s_{1}}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{q_{2}}}\|g_{k}\|_{L_{v}^{p_{2}}},\quad 1\leq p_{2}<2.$
Estimate for $Q_{1}^{s_{2}}$
Using (4.19) and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k}|Q_{1}^{s_{2}}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}$
@@
$\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f_{l}(v^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q_{1}}}\right)\left(\frac{{\langle}v{\rangle}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q_{1}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<|v|^{-1/2}}\,d\sigma\,dv_{1}$
(4.24) $\lesssim I_{1}^{1/p_{1}}(v)I_{2}^{1/q_{1}}(v),$
where
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{p_{1}}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p_{1}/q_{1}}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1},$
@@
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{q_{1}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq_{1}}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Since $l>2$, (2.26) yields
$I_{2}\lesssim{\langle}v{\rangle}^{q_{1}-2}\|g_{k}\|_{L_{v}^{q_{1}}}^{q_{1}}.$
Notice again that the right-hand side of this estimate grows as $|v|\to\infty$ since $q_{1}>2$. However, we will see that the estimate for $I_{1}$ provides sufficient decay to counteract this growth.
Indeed, in order to estimate $I_{1}$, we use the same argument as for $Q_{1}^{s_{1}}$ using the definition (4.16) of $\mathcal{B}_{2}(v)$ instead (which imposes $|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<{\langle}v{\rangle}^{-1/2}$) and exchanging the exponents. We obtain
@@
$I_{1}(v)\lesssim\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\mathds{1}_{|v|>10}\int_{0}^{{\langle}v{\rangle}^{-1/2}}x^{1-\frac{2p_{1}}{q_{1}}}\,dx\,d\nu$ $\lesssim{\langle}v{\rangle}^{\frac{p_{1}}{q_{1}}-1}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\,d\nu$
$={\langle}v{\rangle}^{\frac{p_{1}}{q_{1}}-1}\|f_{l}\|_{L_{v}^{p_{1}}}^{p_{1}},$
where we used the fact that $p_{1}<2<q_{1}$ for the convergence of the integral in $x$.
Combining (4.24) with the bounds for $I_{1},I_{2}$, and using the exponents’ condition, we obtain
${\langle}v{\rangle}^{k}|Q_{1}^{s_{2}}(f,g)(v)|$ $\lesssim|v|^{\frac{1}{q_{1}}-\frac{1}{p_{1}}+1-\frac{2}{q_{1}}}\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}}=\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}},$
therefore
(4.25) $\|Q_{1}^{s_{2}}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}}.$
@@ 4.3. The $L_{v}^{r}$– estimate for $Q_{0}$
Let $0<\gamma\leq 1$ and $2\leq r<\infty$. Consider $1\leq p_{0}\leq 2\leq q_{0}\leq r$ with $\frac{1}{p_{0}}+\frac{1}{q_{0}}=1+\frac{1}{r}$. Then, for any $k\geq l>2$ there holds the estimate
(4.26) $\|{\langle}v{\rangle}^{k}Q_{0}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L_{v}^{p_{0}}}\|{\langle}v{\rangle}^{l}g\|_{L_{v}^{q_{0}}},\quad\forall\,f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
@@
Let $k\geq l>2$ and fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Denote $f_{k}:={\langle}v{\rangle}^{k}f$, $g_{l}:={\langle}v{\rangle}^{l}g$. We note that, since $p_{0}\leq 2\leq q_{0}$, we have $q_{0}^{\prime}\leq 2\leq p_{0}^{\prime}$ where $p_{0}^{\prime},q_{0}^{\prime}$ are the conjugate exponents of $p_{0},q_{0}$ respectively.
We use the decomposition (4.5) from Proposition 4.1.
Estimate for $Q_{0}^{sm},Q_{0}^{r}$
The operators $Q_{0}^{sm}$, $Q_{0}^{r}$ can be treated similarly to Proposition 4.1, we just use Theorem 3.2 instead of Theorem 3.1. Omitting details, we obtain
(4.27) $\|{\langle}v{\rangle}^{k}Q_{0}^{sm}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{0}}}\|g_{k}\|_{L_{v}^{q_{0}}},$
(4.28) $\|{\langle}v{\rangle}^{k}Q_{0}^{r}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{0}}}\|g_{k}\|_{L_{v}^{q_{0}}}.$
Estimate for $Q_{0}^{s}$
We take different cases for $1<p_{0}\leq 2\leq q_{0}<r$ and $p_{0}=1,\,q_{0}=r$.
$\bullet$ Case $1<p_{0}\leq 2\leq q_{0}<r$ : Using (4.8) and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}|f_{k}(v^{*})g_{l}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}|f_{k}(v^{*})|^{p_{0}/q_{0}^{\prime}}}{{\langle}v_{1}^{*}{\rangle}^{l}}\right)\left(|g_{l}(v_{1}^{*})|^{q_{0}/p_{0}^{\prime}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}\right)\left(|f_{k}(v^{*})|^{p_{0}/r}|g_{l}(v_{1}^{*})|^{q_{0}/r}\right)$
$\hskip 142.26378pt\times\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.29) $\leq I_{1}^{1/q_{0}^{\prime}}(v)I_{2}^{1/p_{0}^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})|^{p_{0}}|g_{l}(v_{1}^{*})|^{q_{0}}\,d\sigma\,dv_{1}\right)^{1/r},$

### v4

@@ 4.1. The $L_{v}^{\infty}$– estimate
$J_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq}}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}.$
Since $|v_{1}^{*}|^{2}>\frac{E}{2}>\frac{|v|^{2}}{2}>8$ and $ql>pl>2$, we use (2.26) to bound
@@
$J_{2}(v)\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq}}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}\lesssim\|g_{k}\|_{L^{q}_{v}}^{q}.$
In order to estimate $J_{1}$, we use a symmetric argument as for $I_{1}$. Namely, we use the definition (4.4) of $\mathcal{B}_{2}(v)$, followed by (2.6), (2.13) and the substitution $y:=u=v-v_{1}$ to write
@@
$J_{1}(v)$ $\leq{\langle}v{\rangle}^{1-\frac{p}{q}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(u))|^{p}\mathds{1}_{|v|>4}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{p(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{q/p}}\,dv_{1}\,d\sigma$
@@
$={\langle}v{\rangle}^{1-\frac{p}{q}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(y))|^{p}\mathds{1}_{|v|>4}\frac{|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{p(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{q/p}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ on the right-hand side of the above bound, as well as the fact that $p(k-\gamma)\geq 2$, to obtain
@@
$J_{1}(v)$ $\lesssim{\langle}v{\rangle}^{1-\frac{p}{q}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p}\mathds{1}_{|v|>4}\frac{|\widehat{\nu}\cdot\sigma|^{p(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{p/q}}\,d\nu\,d\sigma$
$\leq{\langle}v{\rangle}^{1-\frac{p}{q}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p}\mathds{1}_{|v|>4}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{p/q}}\,d\sigma\,d\nu$
$\approx{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}<{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p/q}}\,dx\,d\nu.$
The inner integral now can be estimated as follows
$\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}<{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p/q}}\,dx\lesssim\int_{0}^{1}\frac{\mathds{1}_{1-x<{\langle}v{\rangle}^{-1}}}{(1-x)^{p/q}}\approx\int_{0}^{{\langle}v{\rangle}^{-1}}z^{-p/q}\,dz\lesssim{\langle}v{\rangle}^{\frac{p}{q}-1},$
where we used the fact that $p<q$ for the convergence of the integral in $z$. We obtain
$J_{1}(v)\lesssim\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p}\,d\nu=\|f_{l}\|_{L^{p}}^{p}.$
Combining (4.10) with the bounds for $J_{1},J_{2}$, we obtain
(4.11) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}_{s,2}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}.$
@@ 4.2. The $L_{v}^{r}$– estimate
Let $0<\gamma\leq 1$ and $b$ satisfying (H.1)–(H.3). Consider $2<r<\infty$, $1<p<2<q<\infty$ with $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$ and $k\geq l>1$ with $k\geq 2+\gamma-\frac{2}{q}$, $l>2-\frac{2}{q}$. Then, there holds the estimate
(4.12) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},\quad\forall\,f,g\in C_{c}({\mathbb{R}}^{3}).$
@@
Let $k\geq l>1$ with $k\geq 2+\gamma-\frac{2}{q}$, $l>2-\frac{2}{q}$. Note that these assumptions imply $q^{\prime}(k-\gamma)\geq 2$ and $q^{\prime}l>2$. Consider $f,g\in C_{c}({\mathbb{R}}^{3})$, and denote again $f_{l}:={\langle}v{\rangle}^{l}f$, $g_{k}:={\langle}v{\rangle}^{k}g$.
We recall the decomposition (4.2) from Proposition 4.1.
Estimate for $Q_{sm}^{+}$, $Q_{r}^{+}$
The operators $Q_{sm}^{+},Q_{r}^{+}$ can be treated similarly to Proposition 4.1 using Theorem 3.2 instead of Theorem 3.1. Omitting the details, we obtain
(4.13) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{sm}^{+}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}+\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}},$
(4.14) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{r}^{+}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}+\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}}.$
We now estimate the singular operators $Q_{s,1}^{+}$ and $Q_{s,2}^{+}$.
Estimate for $Q_{s,1}^{+}$
Since $q>2$, we have $\frac{1}{r}<\frac{1}{p}-\frac{1}{q}=\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}$. Hence we can choose $\alpha>1$ such that $\frac{\alpha}{r}<\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}$. Relying on the exponents’ compatibility condition, we can write
(4.15) $1=\frac{1}{p^{\prime}}+\frac{1}{q^{\prime}}+\frac{1}{r}=\left(\frac{1}{p^{\prime}}-\frac{1}{q^{\prime}}+\frac{\alpha}{r}\right)+\frac{2}{q^{\prime}}+\frac{1-\alpha}{r}.$
We argue by duality; we fix $\phi\in L_{v}^{r^{\prime}}$, and estimate the quantity
$I(\phi):=\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{k+1-\gamma}Q_{s,1}^{+}(v)\phi(v)\,dv.$
Using (4.7), the fact that $|v|\lesssim|v_{1}^{*}|$, (4.15) and Hölder’s inequality, we bound
$|I(\phi)|$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}|\phi(v)|\,d\sigma\,dv_{1}\,dv$
$\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v_{1}^{*}{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}|\phi(v)|\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v_{1}^{*}{\rangle}^{\frac{1}{p^{\prime}}-\frac{1}{q^{\prime}}+\frac{\alpha}{r}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-1}|f_{l}(v^{*})|^{q/p^{\prime}}|\phi(v)|^{r^{\prime}/p^{\prime}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}\right)$
$\hskip 85.35826pt\times\left(\frac{{\langle}v_{1}^{*}{\rangle}^{2/q^{\prime}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q^{\prime}}|g_{k}(v_{1}^{*})|^{p/q^{\prime}}|\phi(v)|^{r^{\prime}/q^{\prime}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 85.35826pt\times\left(\frac{{\langle}v_{1}^{*}{\rangle}^{\frac{1-\alpha}{r}}|f_{l}(v^{*})|^{q/r}|g_{k}(v_{1}^{*})|^{p/r}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha/r}}\right)\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}}\,d\sigma\,dv_{1}\,dv$
(4.16) $\leq I_{1}^{1/p^{\prime}}I_{2}^{1/q^{\prime}}I_{3}^{1/r},$